OpenStudy (anonymous):

if the slope of the curve x^2 +y^2 =25 is equal to -(3/4), what is the point of tangency? A.(2,3) b.(3,4) c.(3,4,2) d.(3,5,4)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

last 2 answers don't even make sense!

OpenStudy (callisto):

diff both sides wrt x, 2y(dy/dx) +2x =0 =>dy/dx=-2x/2y = -x/y the sub number to get -3/4 that is B

OpenStudy (anonymous):

thx for the help guys